# O. Al-Ghattas, J. Chen, D. Sanz-Alonso

### v1

@@ 5 Conclusions, discussion, and future directions
This paper has established sharp dimension-free concentration inequalities for simple random tensors. In so doing, we have developed the theory of multi-product empirical processes, leveraging generic chaining methods to obtain sharp high-probability upper bounds on their suprema. Our framework extends classical results on quadratic and product empirical processes to higher-order settings.
@@
In this paper, we focused on Gaussian and sub-Gaussian random variables in developing tensor concentration inequalities, as well as on sub-Gaussian function classes $\mathcal{F}$ in the study of multi-product empirical processes. An important direction for future research is to extend these results to more general measures and weakly bounded empirical processes, such as those associated with isotropic log-concave measures [23, 47, 48, 51] and Gaussian mixtures [6, Section 4]. We anticipate that our analysis can be extended to general distributions by further incorporating the powerful framework developed in [51]. As shown in [51, Section 4.3], the graded version of $\gamma$-type functionals offers advantages over the $\psi_{2}$-based complexity measure when analyzing unconditional log-concave ensembles. We expect these favorable properties to persist as we extend the results to higher-order settings.
@@
[…]
@@
()${\leq}\sqrt{\log_{2}(1+x)}+\sqrt{\log_{2}(1+y)}+\sqrt{\frac{\left(\log_{2}(1+x)+\log_{2}(1+y)\right)^{m}}{N}}$
@@
[…]

### v2

@@ 5 Conclusions, discussion, and future directions
This paper has established sharp dimension-free concentration inequalities for simple random tensors. In so doing, we have developed the theory of $L_{p}$ empirical processes, leveraging generic chaining methods to obtain sharp high-probability upper bounds on their suprema. Our framework extends classical results on quadratic and product empirical processes to higher-order settings.
@@
In this paper, we focused on Gaussian and sub-Gaussian random variables in developing tensor concentration inequalities, as well as on sub-Gaussian function classes $\mathcal{F}$ in the study of $L_{p}$ empirical processes. An important direction for future research is to extend these results to more general measures and weakly bounded empirical processes, such as those associated with isotropic log-concave measures [23, 47, 48, 51] and Gaussian mixtures [6, Section 4]. We anticipate that our analysis can be extended to general distributions by further incorporating the powerful framework developed in [51]. As shown in [51, Section 4.3], the graded version of $\gamma$-type functionals offers advantages over the $\psi_{2}$-based complexity measure when analyzing unconditional log-concave ensembles. We expect these favorable properties to persist as we extend the results to higher-order settings.
@@
$\overset{\text{($\star\star$)}}{=}\inf_{g\in\mathcal{F}}\sup_{f\in\mathcal{F}}\|f-g\|_{\psi_{2}}\overset{\text{($\star\star\star$)}}{\geq}\frac{1}{2}\text{diam}_{\psi_{2}}(\mathcal{F}),$
where ($\star$) follows by the positivity of the terms in the sum for $s\geq 1$, ($\star\star$) holds by the fact that the infimum is over all admissible sequences with $|\mathcal{F}_{0}|=1$, and ($\star\star\star$) holds by a standard bound on $\inf_{g\in\mathcal{F}}\sup_{f\in\mathcal{F}}\|f-g\|_{\psi_{2}}$, which is referred to in the literature as the Chebyshev radius of $\mathcal{F}$. Now, in the symmetric case, we have
@@
$\overset{\text{($\star$)}}{\leq}\sqrt{\log_{2}(1+x)}+\sqrt{\log_{2}(1+y)}+\sqrt{\frac{\left(\log_{2}(1+x)+\log_{2}(1+y)\right)^{m}}{N}}$
@@
$\mathbb{E}\varphi\bigg(\frac{|\sum_{i=1}^{N}Y_{i}-\mathbb{E}Y_{i}|}{c(m)\sqrt{N}M}\bigg)$ $=\int_{0}^{\infty}\mathbb{P}\bigg(\varphi\bigg(\frac{|\sum_{i=1}^{N}Y_{i}-\mathbb{E}Y_{i}|}{c(m)\sqrt{N}M}\bigg)\geq u\bigg)du$
$=\int_{0}^{\infty}\mathbb{P}\bigg(\frac{|\sum_{i=1}^{N}Y_{i}-\mathbb{E}Y_{i}|}{c(m)\sqrt{N}M}\geq\varphi^{-1}(u)\bigg)du$
$\leq 1+\int_{1}^{\infty}\mathbb{P}\bigg(\frac{|\sum_{i=1}^{N}Y_{i}-\mathbb{E}Y_{i}|}{c(m)\sqrt{N}M}\geq\varphi^{-1}(u)\bigg)du$